# The Effect of Compression Techniques on Large Multimodal Language Models in the Medical Domain

### v1

@@ Abstract
[…]
@@ 3.2 Structural Pruning
We store the outputs from the forward passes. We then compare the outputs using cosine similarity. To compute the cosine similarity we utilize sentence transformer [16] embedding model $all-MiniLM-L6-v2$. We denote the sentence transformer model as $g(\cdot)$. The embeddings are calculated as
@@ 3.4 Post Training Quantization
[…]
@@
[…]
@@ 4 Experiments
Our experiments were done using Kaggle’s T4 GPUs. Our codes are stored in Github.
@@ 4.2 Evaluation Metrics
To evaluate performance, we use LLM as a judge for visual question answering. We denote the score obtained from Judge LLM as the performance score. The details of the implementation is described in Appendix B. To evaluate the efficiency of our model, we use the Peak VRAM usage and latency as our metrics.
@@ 4.3 Structural Pruning
Structural pruning removes parameters from the model. While it reduces the number of parameters providing faster inference and less VRAM usage, it also degrades the model’s conversation capability to some extent. After removing each block, we obtain an importance score for each layer in the model. A higher score indicates better output accuracy even without the corresponding layer, signifying the lower importance of that block. Then, we removed entire layers from the model based on the importance scores. Layers with minimum sum were identified and pruned iteratively. The results are shown in Table 1 which illustrates the tradeoff between model performance and efficiency.
@@ 4.4 Supervised Finetuning
After structured pruning,we perform supervised finetuning on our model. Table 2 shows the effect of pruning on finetuning memory requirement and time. We use Qlora finetuning method with $rank=16$ and $alpha=8$.
@@ 4.6.1 Performance and Efficiency Comparison
Table 4 shows the overall result for our methodology and our baselines.
@@ 4.6.2 Comparison
We compare our method to traditional implementations to understand the impact of each component in our pipeline. Table 5 shows the impact of removing each components of our pipeline.
@@
Bitsandbytes quantization 47 --7% accuracy drop
Pruning Based on Magnitude Baseline 50 --4% accuracy drop
@@ 4.7 Case Studies
Figure 5 presents qualitative comparisons between full-precision and compressed models. Our method maintains diagnostic fidelity in most cases, with minor degradation in ambiguous queries.

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) hold huge potential for usage in the medical domain, but their computational costs necessitate efficient compression techniques. This paper evaluates the impact of structural pruning and activation-aware quantization on a fine-tuned LLAVA model for medical applications. We propose a novel layer selection method for pruning, analyze different quantization techniques, and assess the performance trade-offs in a prune-SFT-quantize pipeline. Our proposed method enables MLLMs with 7B parameters to run within 4 GB of VRAM, reducing memory usage by $70\%$ while achieving $4\%$ higher model performance compared to traditional pruning and quantization techniques in the same compression ratio.
@@ 3.2 Structural Pruning
We store the outputs from the forward passes. We then compare the outputs using cosine similarity. To compute the cosine similarity we utilize sentence transformer [16] embedding model allMiniLML6v2all-MiniLM-L6-v2. We denote the sentence transformer model as $g(\cdot)$. The embeddings are calculated as
@@ 3.4 Post Training Quantization
Figure 4 illustrates the distribution of activation values at the output of the query projection (q_proj) matrix in the 30th transformer layer of the language model. The input tensor is projected using a query weight matrix of shape $4096\times 4096$, resulting in an output activation matrix of shape $\text{input\_size}\times 4096$. While the majority of activation values lie within the range of 0 to 3, several prominent outliers exceed a value of 10, with the highest peak surpassing 25. These outliers suggest the presence of sparsely distributed, high-magnitude activations, which may play a critical role in the model’s attention mechanism and information propagation.
@@
$s=s_{X}^{\alpha},\quad\alpha^{*}=\underset{\alpha}{\arg\min}\;\mathcal{L}(s_{X}^{\alpha}).$ (8)
@@ 4 Experiments
This section details the experimental setup designed to evaluate our proposed compression pipeline. We assess the trade-offs between model performance and computational efficiency (VRAM usage, latency) at each stage. All experiments were conducted on Kaggle’s T4 GPUs, and the source code is available on Github.
@@ 4.2 Evaluation Metrics
To assess model quality in the visual question answering task, we employ an LLM-based judge (detailed in Appendix B) to generate a performance score. This approach was chosen over traditional metrics to better capture the nuance and clinical relevance of the generated text. For computational efficiency, we measure peak VRAM usage (GB) and inference latency ( ms/token ).
@@ 4.3 Structural Pruning
Structural pruning removes parameters from the model. While it reduces the number of parameters providing faster inference and less VRAM usage, it also degrades the model’s conversation capability to some extent. After removing each block, we obtain an importance score for each layer in the model. A higher score indicates better output accuracy even without the corresponding layer, signifying the lower importance of that block. Then, we removed entire layers from the model based on the importance scores. Layers with minimum sum were identified and pruned iteratively. Layers with the lowest importance scores were then iteratively removed. The results, presented in Table 1, illustrate the trade-off between model performance and efficiency. As the table shows, performance remains relatively stable with up to four layers removed (a 65.75 score at 10% compression), but degrades sharply thereafter. Removing ten layers (29% compression) causes the performance score to collapse to 27.25, indicating that crucial model capabilities have been lost.
@@ 4.4 Supervised Finetuning
To counteract the performance degradation caused by pruning, we applied supervised finetuning (SFT) to the pruned models using the Qlora method (rank = 16, alpha = 8). Table 2 demonstrates that SFT is highly effective at recovering performance. For example, the 5.6B parameter model’s score improved from 62.50 to 74.25 after finetuning. The table also confirms that finetuning a smaller model is more efficient, requiring less VRAM and time.
@@ 4.6.1 Performance and Efficiency Comparison
Table 4 consolidates the results, comparing our complete pipeline against the baselines. The findings highlight the effectiveness of our combined approach. Our final model (Prune + SFT + Quant) achieves a performance score of 54.25 while requiring only 3.9 GB of VRAM. This is a dramatic improvement over a naive ’Prune + Quant’ approach, which scores an unusable 25.00. While performance is lower than the 13.4 GB original model, our method successfully reduces VRAM usage by 71%, achieving a strong balance between performance and efficiency.
@@ 4.6.2 Comparison
To isolate the benefits of our specific component choices, we performed an ablation study detailed in Table 5. The results validate our methodology: replacing our activation-aware quantization (AWQ) with a standard Bitsandbytes technique causes a 7-point performance drop. Furthermore, using a simpler magnitude-based pruning baseline leads to a 4-point drop compared to our method. Attempting to use a standard round-to-neighbor quantization method resulted in unintelligible output, underscoring the necessity of our pipeline’s advanced techniques.
@@
Bitsandbytes quantization 47 -7% accuracy drop
Pruning Based on Magnitude Baseline 50 -4% accuracy drop
@@ 4.7 Case Studies
Figure 5 provides a qualitative comparison of model responses at different compression stages. The Original Model gives a detailed, accurate clinical description. After naive pruning and quantization (Pruned + Quantized), the model’s response becomes generic and less clinically useful. However, the output from our full pipeline (Pruned + SFT + Quantized) recovers significant diagnostic detail, correctly identifying key characteristics of the condition. This case study demonstrates that our method preserves essential domain-specific knowledge even after significant compression.